OpenStudy (anonymous):

hey guys i need some clarification regarding the integration and differentiation of natural log and exponential functions...help?

OpenStudy (anonymous):

like for example: how would you find the integral of \[x^{2x}\]

OpenStudy (anonymous):

would it follow the integration rule of k^x where the integral is k^x/ln k times dx

OpenStudy (anonymous):

general rules of integrals and derivatives of log and natural log would be very helpful

OpenStudy (anonymous):

No. k^x formula is used ONLY when the constant is given, i.e. 3^x of 4^(2x).

OpenStudy (anonymous):

yes k is a constant

OpenStudy (anonymous):

But x^(2x) is NOT of the form (a^x), where a is a constant.

OpenStudy (anonymous):

so which integral rule would be used to solve this?

OpenStudy (anonymous):

when would i use the chain rule and when would i not?

zepdrix (zepdrix):

Hmm I don't think a solution exists to such a problem.. Was it in your book or something..? :o Example:\[\Large y\quad=\quad x^{2x}\]Taking the log of each side gives us,\[\Large \ln y\quad=\quad 2x\ln x\]Integrating each side with respect to it's proper variable gives us,\[\Large y \ln y- y\quad=\quad x^2 \ln x-\frac{1}{2}x+C\]And then we could ummmmm maybe... hmm

OpenStudy (anonymous):

ummmmm is right you are not going to find a nice closed form for this function \[x^{2x}=e^{2x\ln(x)}\] and you can pretty much forget finding an anti derivative for it you can find the derivative no problem, but not an anti derivative

OpenStudy (anonymous):

yes its in my math book

OpenStudy (anonymous):

i was actually just asking for general rules regarding the use of chain rules when integrating or find the derivative of the natural log and exponential functions...

zepdrix (zepdrix):

Like these?\[\Large \int\limits \ln x, \qquad\qquad \int\limits a^x\;dx\]

OpenStudy (anonymous):

because sometimes, as seen from my book, you would multiply the derivative of the inner function to ur outer functions and sometimes you wouldn't...

OpenStudy (anonymous):

yea how those equal to 1/x and \[(a ^{x})/\ln a\]

OpenStudy (anonymous):

respectively

OpenStudy (anonymous):

and this is where my problem lies

OpenStudy (anonymous):

for \[\int\limits_{?}^{?}a ^{x}dx\]

OpenStudy (anonymous):

would it be \[(a ^{x})/\ln a\] and then multiplied by the derivative of the inner function x?

OpenStudy (anonymous):

or just \[(a ^{x})/\]?

zepdrix (zepdrix):

There isn't a chain rule for integration. At least I wouldn't describe it that way. So you remember this derivative, yes?\[\Large (a^x)'\quad=\quad a^x \ln a\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and integral would be divide by ln a

zepdrix (zepdrix):

So if we were to take this integral,\[\Large \int\limits \color{orangered}{a^x \ln a}\; dx\]What would we expect to get?

zepdrix (zepdrix):

Yah good, it gives us back what we started with in that example.

zepdrix (zepdrix):

I'm not sure what you mean by `multiply by the inner function x`. Like if we had:\[\Large \int\limits a^{2x}\;dx\]Then yes, we get a little bit something more than just dividing by the ln(a). Is that what you're asking about?

OpenStudy (anonymous):

are you sure there isn't a chain rule for integration? i thought it applies to everything

OpenStudy (anonymous):

yes so would u multiply it by 2 (the derivative of the inner function: 2x)? or not?

zepdrix (zepdrix):

The `chain rule` (taking the derivative of a^2x) would produce an extra factor of 2. So when we integrate, we need to divide by 2 to compensate for the extra factor of 2.

zepdrix (zepdrix):

\[\Large (e^{2x})'\quad=\quad 2e^{2x}\]We ended up with a factor of 2 when differentiating, So when we integrate, we need to divide by 2,\[\Large \int\limits e^{2x}dx\quad=\quad \frac{1}{2}e^{2x}+C\]

OpenStudy (anonymous):

u forgot the 2 for the integral of e^2x that should've cancelled out the 1/2 on the right...

zepdrix (zepdrix):

I was showing the difference between `differentiating e^2x` and `integrating e^2x` but yes you have the right idea :)

OpenStudy (anonymous):

so the chain rule apply to which? sry i suck at this :P

zepdrix (zepdrix):

This works quite nicely when our `inner function` is linear. If it's linear with a coefficient, we simply divide by that coefficient when we integrate. Unfortunately, unlike with taking a derivative, this trick doesn't always work. If the inner function is NOT linear, we have something else going on. Example:\[\Large \left(e^{x^2}\right)'\quad=\quad (2x)e^{x^2}\]So taking a derivative produces a 2x. Integrating the same function:\[\Large \int\limits e^{x^2}\;dx\]does not allow us to simply divide by the derivative of the inner function (2x).

zepdrix (zepdrix):

Mmm maybe I'm making it more complicated than it needs to be.. hmm thinking..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

actually

OpenStudy (anonymous):

for integration u would divide it by the derivative of the inner function

OpenStudy (anonymous):